OpenStudy (anonymous):

Does anybody know how to integrate: ? 0 to 50 ʃ 1- (t/100)^1.5 dt

OpenStudy (anonymous):



OpenStudy (agreene):

\[\int (1-(\frac{t}{100})^{1.5})dt=t-0.0004t^{2.5}\] evaluate across limits.

OpenStudy (anonymous):

can u somehow show me the steps please..will be grateful

OpenStudy (agreene):

\[\int (1-(\frac{t}{100})^{1.5})dt = \int (1-0.001t^{1.5})dt=\int 1 dt-\int 0.001t^{1.5}dt\] \[=\int 1 dt - 0.001\int t^{1.5}dt=t-0.0004t^{2.5}\]

OpenStudy (anonymous):

Thanks